# Reinforcement Learning for a Discrete-Time Linear-Quadratic Control Problem with an Application

### v1

@@ 2 Formulation
$V^{u}(0,x_{0},y_{0})=\min_{u}~{}\mathbb{E}\left[\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}Q_{T}\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}\right].$ (2)
@@
$J(t,x_{t},y_{t})$ $=\mathbb{E}\left[\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}Q_{T}\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}+\lambda\sum_{t=0}^{T-1}\int_{\mathbb{R}^{m}}\pi_{t}(u)\ln\pi_{t}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\right],$ (4)
@@
$=\min_{\pi_{t},\cdots,\pi_{T-1}}\mathbb{E}\left[\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}Q_{T}\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}+\lambda\sum_{s=t}^{T-1}\int_{\mathbb{R}^{m}}\pi_{s}(u)\ln\pi_{s}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\right],$
@@ 3 Mean–Variance Model of Asset-liability Management
$\left\{\begin{array}[]{rl}\displaystyle\min_{u}&\mbox{Var}\left(X_{T}^{u}-l_{T}\right)\equiv\mathbb{E}[(X_{T}-l_{T})^{2}]-d^{2},\\[5.69054pt] \mbox{s.t.}&\mathbb{E}\left[X_{T}^{u}-l_{T}\right]=d,\\[5.69054pt] \mbox{s.t.}&X_{t+1}^{u}=r_{f}X_{t}^{u}+r_{t}u_{t},\\[5.69054pt] &l_{t+1}=q_{t}l_{t},\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mbox{for }t=0,1,\cdots,T-1,\end{array}\right.$ (9)
@@
$\min_{u}~{}\mathbb{E}[(X_{T}-l_{T})^{2}]-d^{2}-2\gamma\left(\mathbb{E}[X_{T}-l_{T}]-d\right)^{2}=\min_{u}~{}\mathbb{E}[(X_{T}-\gamma-l_{T})^{2}]-(\gamma-d)^{2}.$ (10)
@@
$J(t,x_{t},l_{t})$ $=\mathbb{E}\left[\begin{pmatrix}x_{T}\\ l_{T}\end{pmatrix}Q_{T}\begin{pmatrix}x_{T}\\ l_{T}\end{pmatrix}+\lambda\sum_{t=0}^{T-1}\int_{\mathbb{R}^{m}}\pi_{t}(u)\ln\pi_{t}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\right],$ (13)
@@ 4.1 Policy improvement and policy convergence
$J^{\pi^{j}}(t,x_{t},l_{t})$ $=\begin{pmatrix}x_{t}\\ l_{t}\end{pmatrix}M_{t+j}\begin{pmatrix}x_{t}\\ l_{t}\end{pmatrix}+\frac{\lambda}{2}j\ln\bigg{(}\frac{B^{2}+D^{2}}{\pi\lambda}\bigg{)}+\frac{\lambda}{4}(j-1)j\ln\left(\frac{A^{2}D^{2}}{B^{2}+D^{2}}\right)$
@@ 4.2 Algorithm Design
Then we need to learn $\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5}$
@@
$J^{\pi}(t,x_{t},y_{t})=\mathbb{E}\bigg{[}J^{\pi}(t+1,x_{t+1},y_{t+1})+\lambda\int_{\mathbb{R}}\pi_{t}(u)\ln\pi_{t}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\bigg{]}.$
Re-arranging equation (4.2), we obtain

### v2

@@ 2 Formulation
$V^{u}(0,x_{0},y_{0})=\min_{u}~{}\mathbb{E}\left[\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}^{\prime}Q_{T}\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}\right].$ (2)
@@
$J(t,x_{t},y_{t})$ $=\mathbb{E}\left[\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}^{\prime}Q_{T}\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}+\lambda\sum_{t=0}^{T-1}\int_{\mathbb{R}^{m}}\pi_{t}(u)\ln\pi_{t}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\right],$ (4)
@@
$=\min_{\pi_{t},\cdots,\pi_{T-1}}\mathbb{E}\left[\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}^{\prime}Q_{T}\begin{pmatrix}x_{T}\\ y_{T}\end{pmatrix}+\lambda\sum_{s=t}^{T-1}\int_{\mathbb{R}^{m}}\pi_{s}(u)\ln\pi_{s}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\right],$
@@ 3 Mean–Variance Model of Asset-liability Management
$\left\{\begin{array}[]{rl}\displaystyle\min_{u}&\mbox{Var}\left(X_{T}^{u}-l_{T}\right)\equiv\mathbb{E}[(X_{T}^{u}-l_{T})^{2}]-d^{2},\\[5.69054pt] \mbox{s.t.}&\mathbb{E}\left[X_{T}^{u}-l_{T}\right]=d,\\[5.69054pt] &X_{t+1}^{u}=r_{f}X_{t}^{u}+r_{t}u_{t},\\[5.69054pt] &l_{t+1}=q_{t}l_{t},\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mbox{for }t=0,1,\cdots,T-1,\end{array}\right.$ (9)
@@
$\min_{u}~{}\mathbb{E}[(X_{T}^{u}-l_{T})^{2}]-d^{2}-2\gamma\left(\mathbb{E}[X_{T}^{u}-l_{T}]-d\right)=\min_{u}~{}\mathbb{E}[(X_{T}^{u}-\gamma-l_{T})^{2}]-(\gamma-d)^{2}.$ (10)
@@
$J(t,x_{t},l_{t})$ $=\mathbb{E}\left[\begin{pmatrix}x_{T}\\ l_{T}\end{pmatrix}^{\prime}Q_{T}\begin{pmatrix}x_{T}\\ l_{T}\end{pmatrix}+\lambda\sum_{t=0}^{T-1}\int_{\mathbb{R}^{m}}\pi_{t}(u)\ln\pi_{t}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\right],$ (13)
@@ 4.1 Policy improvement and policy convergence
$J^{\pi^{j}}(t,x_{t},l_{t})$ $=\begin{pmatrix}x_{t}\\ l_{t}\end{pmatrix}^{\prime}M_{t+j}\begin{pmatrix}x_{t}\\ l_{t}\end{pmatrix}+\frac{\lambda}{2}j\ln\bigg{(}\frac{B^{2}+D^{2}}{\pi\lambda}\bigg{)}+\frac{\lambda}{4}(j-1)j\ln\left(\frac{A^{2}D^{2}}{B^{2}+D^{2}}\right)$
@@ 4.2 Algorithm Design
Then we need to learn $\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5}$.
@@
$J^{\pi}(t,x_{t},y_{t})=\mathbb{E}\bigg{[}J^{\pi}(t+1,x_{t+1},y_{t+1})+\lambda\int_{\mathbb{R}}\pi_{t}(u)\ln\pi_{t}(u)\mathrm{d}u\bigg{|}\mathcal{F}_{t}\bigg{]}.$ (18)
Re-arranging equation (18), we obtain